OpenStudy (anonymous):

How do I find the surface area of a cylinder with base radius of 3 ft and height of 8ft with the answer in terms of pi?

OpenStudy (kirbykirby):

You can think of a cylinder being a rectangle rolled up, with two circles covering the 2 bases. Thus, the surface area will be the are of that rectangle, plus the area of the 2 circles. The dimensions of the rectangle: the length is the height of the cylinder the width is the circumference of the circle So the area of the rectangle = (length)*(width) = (height)*(circumference) And the area of one circle is: \(\pi r^2\), where \(r\) is the radius.

OpenStudy (anonymous):

use the formula: A=2πrh+2πr^2 A=2π(3)(8)+2π(3)2 A=207.35 sq feet

OpenStudy (triciaal):

with the answer in terms of pi? means do not substitute the value of pi

OpenStudy (triciaal):

@kirbykirby good job at deriving the formula @gregory.d A = 2*pi*r*h + 2*pi*r^2 as above = 2*pi*r(h + r) substituting for r and h 2*pi*3 ( 3 + 8) = 66 pi